# Higher-Order Staircase Codes

### v1

@@ IV-F Notions of Optimality for $(L,M)$-DTS Constructions
As with Golomb rulers which correspond to $(1,M)$-DTSs, it is easy to construct an $(L,M)$-DTS for arbitrary $L$ and $M$ if no further objectives are placed. A well-studied problem is that of constructing an $(L,M)$-DTS of minimum scope where scope is the largest mark across all rulers when rulers are assumed to be translated so that $d_{0}^{(\ell)}=0$ for all $\ell\in[L]$. Under our notation, the scope of an $(L,M)$-DTS is simply $\max_{\ell\in[L]}d_{M}^{(\ell)}$, the largest ruler length. Observe next that we have a trivial lower bound for any $(L,M)$-DTS that
@@
with equality if and only if the union of the sets of differences of the $L$ rulers is precisely $\{1,2,\dots,L(M+1)M/2\}$.
An $(L,M)$-DTS with minimum scope, or equivalently, minimum maximum length, minimizes decoding memory as per (23). From (23) and (24), we see that the best decoding memory that we can hope to achieve is
$\left(\frac{S}{L}\right)^{2}\left(1+L^{2}\cdot\frac{(M+1)M}{2}\right)=S^{2}\cdot\left(\frac{1}{L^{2}}+\frac{(M+1)M}{2}\right)$
which suggests some decoding memory reduction by increasing $L$.
On the other hand, we see from (22) and (21) that encoding memory is minimized by choice of $(L,M)$-DTS with minimum average length $\frac{1}{L}\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$ rather than minimum maximum length $\max_{\ell\in[L]}d_{M}^{(\ell)}$ which leads to two potentially competing notions of optimality. To the best of the authors’ knowledge, the construction of $(L,M)$-DTSs with minimum average length has not been studied. However, in some special cases, an $(L,M)$-DTS with minimum maximum length also has minimum average length. In particular, we have the following:
@@
An $(L,M)$-DTS for which the scope (or maximum length) lower bound (24) holds with equality also has minimum average length if $M\in\{0,1,2\}$.
@@
In particular, (25) is a sum over all three distances of all $L$ rulers which must be distinct positive integers. This leads the lower bound (26) which is achieved precisely when the scope lower bound (24) is achieved. In the case of $M=1$, the sum of the lengths is precisely the sum of the distances which means that a similar argument holds while the case of $M=0$ is trivial. ∎
Note that if $M>2$, there is no simple relationship such as (25) between a sum over all ruler lengths and the sum over the sums of their respective distances. Moreover, even in the case of $M=2$, a minimum scope $(L,2)$-DTS which does not meet the lower bound (24) does not minimize average length:
@@
Consider the $(2,2)$-DTS of Example 4 given by $\{\{0,2,5\},\{0,6,7\}\}$ which is of scope $7$. Observe that achieving the scope lower bound of $6$ is impossible in this case since it would require equality of (25) which is even and (26) which is $21$. This implies that $\{\{0,2,5\},\{0,6,7\}\}$ has minimum scope. However, the alternative $(2,2)$-DTS given by $\{\{0,3,4\},\{0,2,7\}\}$ which also has minimum scope also has a smaller average length and thus provides an encoding memory reduction if substituted into Example 4.
IV-G Recovery of Tiled Diagonal Zipper codes $(L\geq 1,M=1)$
Taking $M=1$, a minimum scope $(L,1)$-DTS is given by the rulers $\{0,1\},\{0,2\},\dots,\{0,L\}$ which achieve the lower bound (24) and also minimize average length as per Proposition 3. Consider now the higher-order staircase code defined by this $(L,1)$-DTS and the $(2,S/L)$-nets from Example 1 or Example 2. Taking $d_{1}^{(\ell)}=L-\ell$ for all $\ell\in[L]$, we get
@@ IV-G Recovery of Tiled Diagonal Zipper codes $(L\geq 1,M=1)$
belong to some specified component code $\mathcal{C}$ for all $n\in L\mathbb{Z}$ which makes these codes identical to the tiled diagonal zipper codes of [3]. Moreover, by (22), we have an encoding memory of
$\left(\frac{S}{L}\right)^{2}\cdot\frac{L(L+1)}{2}=\frac{1}{2}\left(1+\frac{1}{L}\right)S^{2}$ (27)
which approaches half of the encoding memory of classical staircase codes, corresponding to $(L=1,M=1)$, as $L$ grows.
IV-H An Explicitly Solved Case $(L\geq 1,M=2)$
The case of $M=2$ is the only non-trivial instance of the minimum scope $(L,M)$-DTS problem which has an explicitly constructed solution for arbitrary $L$ [21]. In particular, for all $L\equiv 0\pmod{4}$ and $L\equiv 1\pmod{4}$, an explicitly constructed $(L,2)$-DTS of scope $3L$ is given in [22] and this is the minimum scope since it meets the lower bound (24). Furthermore, for all $L\equiv 2\pmod{4}$ and $L\equiv 3\pmod{4}$, an explicitly constructed $(L,2)$-DTS of scope $13L+$ 1 is given in [23] and it is shown that no smaller scope is possible for such $L$ in [22]. Note, however, that [22, 23] consider a different but equivalent problem to the $(L,2)$-DTS problem so some straightforward translations are needed.
Next, note that when $L\equiv 0\pmod{4}$ or $L\equiv 1\pmod{4}$, and only in these cases, the scope lower bound (24) is met which implies by Proposition 3 that the construction of [22] also minimizes average length. In particular, the minimum average length of $43(3L+1)/$ 4 is also achieved. This yields an encoding memory of
$\left(\frac{S}{L}\right)^{2}\cdot\frac{3L(3L+1)}{4}=\frac{3}{4}\left(3+\frac{1}{L}\right)S^{2}$ (28)
which approaches three-quarters of the encoding memory of generalized staircase codes corresponding to $(L=1,M=2)$ as $L$ grows with $L\equiv 0\pmod{4}$ or $L\equiv 1\pmod{4}$. We emphasize that the construction achieving (28) is explicit and is simply Definition 9 together with the explicit $(L,2)$-DTS of [22]. We only omit the description due to its length.
The encoding memory behaviours (27) and (28) indeed show that our higher-order staircase codes naturally generalize the memory reduction property of the tiled diagonal zipper codes of [3] to higher bit degree.
IV-I The General Case $(L\geq 1,M>2)$
In the case of $M>2$, small scope $(L,M)$-DTSs are typically found by combination of algebraic and computer-search methods [21, 24] and specific values of $L$ and $M$ for which a minimum scope $(L,M)$-DTS is known are documented in [6, Part VI: Ch. 19]. The reader is referred to [6] and the references therein for constructions and construction methods. However, as discussed, we may also be interested in minimizing average length as opposed to scope and this can be beneficial as illustrated by Example 5. We suggest this as a problem for future work.
IV-J Recovery of Self-Orthogonal Convolutional Codes ($S=L$)

### v5

@@ VI Difference Triangle Sets with Minimum Scope and Sum-of-Lengths
for $\ell\in[L]$. As with Golomb rulers or $(1,M)$-DTSs, it is easy to construct an $(L,M)$-DTS for arbitrary $L$ and $M$ if no further objectives are placed. A well-studied hard problem is that of constructing an $(L,M)$-DTS of minimum scope where scope is the largest mark across all (normalized) rulers $\max_{\ell\in[L]}d_{M}^{(\ell)}$. Equivalently, scope is the maximum length among all rulers. We again have a trivial bound on the scope of an $(L,M)$-DTS given by
@@
with equality if and only if the union of the sets of distances of the $L$ rulers is precisely $\{1,2,\dots,L(M+1)M/2\}$. An $(L,M)$-DTS for which the trivial bound holds with equality is said to be perfect. The union of the $L$ (disjoint) sets of distances of the $L$ rulers forming an $(L,M)$-DTS will be referred to concisely as its set of distances.
On the other hand, we also require a minimization of the sum-of-lengths $\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$. To the best of the authors’ knowledge, this problem has not been explicitly studied. We begin by identifying when minimizing scope is equivalent to minimizing sum-of-lengths. Aside from the case of $L=1$ where sum-of-lengths and scope are trivially equivalent, we have the following:
@@
A minimum scope $(L,M)$-DTS also has minimum sum-of-lengths if $M\in\{1,2\}$ and it is perfect.
@@
In particular, (22) is a sum over all three distances of all $L$ rulers which must be distinct positive integers. This leads to the lower bound (23) which is achieved precisely when perfect. When $M=1$, the sum of the lengths is precisely the sum of the distances which means that a similar argument holds. ∎
Beyond this case, the problem of minimizing sum-of-lengths will be distinct and non-trivial. If $M>2$, there is no simple proportionality relationship such as (22) between a sum over all ruler lengths and a sum over all distances. Moreover, even in the case of $M=2$, a minimum scope $(L,2)$-DTS which is not perfect does not necessarily minimize sum-of-lengths as illustrated by the following counterexample:
@@
Consider the $(2,2)$-DTS of Example 4 given by $\{(0,2,5),(0,6,7)\}$ which is of scope $7$. Observe that this scope is already minimum since achieving the scope lower bound of $6$ is impossible since it requires equality of (22) which is even and (23) which is $21$. However, the alternative $(2,2)$-DTS given by $\{(0,3,4),(0,2,7)\}$ which also has minimum scope also has a smaller sum-of-lengths.
Except for the cases of $M=1$ and $M=2$, it remains an open problem to exhibit minimum scope $(L,M)$-DTSs for arbitrary $L$. However, perfect DTSs admit recursive constructions which lead to constructive existence results for infinitely many values of $L$. It is known that perfect DTSs can only exist for $M\leq 4$. Therefore, for the cases of $M=3$ and $M=4$, recursive constructions provide the strongest existence results. The values of $L$ covered by such constructions are often (but not always) too large to be of relevance to applications. Therefore, such constructions do not eliminate the need for computer-search constructions for smaller values of $L$ but are nonetheless at least of theoretical interest. The reader is referred to [34, 35] and the references therein for more on this context. We will study the implications of one such construction due to Wild [36] and Kotzig and Turgeon [37] for the problem of minimizing sum-of-lengths. We will need the following:
Definition 12 (Sharply $2$-transitive permutation group).
A sharply $2$-transitive permutation group $G$ acting on a set $X$ is a group of permutations of $X$ with the following property: For each pair of ordered pairs $(w,x),(y,z)\in X\times X$ with $w\neq x$ and $y\neq z$, there exists a unique permutation $g\in G$ such that $(g(w),g(x))=(y,z)$.
As noted in [36], we can exhibit a finite such group as the group of invertible affine transformations from a finite field $\mathbb{F}_{M+1}$ to itself $x\mapsto ax+b$ where $M+1$ is a prime power and $a,b\in\mathbb{F}_{M+1}$ with $a\neq 0$ yielding $M(M+1)$ such permutations. For $M=3$, the resulting permutations of $[M+1]$ in array format are
$\begin{array}[]{cccccccccccc}0&1&2&3&0&1&2&3&0&1&2&3\\ 1&0&3&2&2&3&0&1&3&2&1&0\\ 2&3&0&1&3&2&1&0&1&0&3&2\\ 3&2&1&0&1&0&3&2&2&3&0&1\\ \end{array}$
and for $M=4$, the permutations are defined by simple modular arithmetic since $\mathbb{F}_{5}$ is $\mathbb{Z}_{5}$.
Theorem 5 (Special case of the combining construction of Wild [36] and Kotzig and Turgeon [37]).
Let $M+1$ be a prime power and let the collection of $M(M+1)$ permutations
$\rho_{k}\colon[M+1]$ $\longrightarrow[M+1]$
$i$ $\longmapsto\rho_{k}(i)$
where $k\in[M(M+1)]$ be a sharply $2$-transitive permutation group acting on $[M+1]$. Let $\mathcal{X}$ be a perfect $(L_{1},M)$-DTS and $\mathcal{Y}$ be a perfect $(L_{2},M)$-DTS with each given respectively by
$\mathcal{X}=\{(x_{0}^{(i)},x_{1}^{(i)},\dots,x_{M}^{(i)})\mid i\in[L_{1}]\}$
and
$\mathcal{Y}=\{(y_{0}^{(i)},y_{1}^{(i)},\dots,y_{M}^{(i)})\mid i\in[L_{2}]\}\text{.}$
Let $\mathcal{W}$ be the set of $L_{1}L_{2}M(M+1)$ rulers given by
$(L_{2}M(M+1)+1)\left(x_{0}^{(i)},x_{1}^{(i)},\dots,x_{M}^{(i)}\right)+\\ \left(y_{\rho_{k}(0)}^{(j)},y_{\rho_{k}(1)}^{(j)},\dots,y_{\rho_{k}(M)}^{(j)}\right)$
for $i\in[L_{1}]$, $j\in[L_{2}]$, and $k\in[M(M+1)]$. Let $\mathcal{Z}$ be the union of $\mathcal{Y}$, $(L_{2}M(M+1)+1)\mathcal{X}$, and $\mathcal{W}$. Then, $\mathcal{Z}$ is a perfect $(L_{1}L_{2}M(M+1)+L_{1}+L_{2},M)$-DTS. (However, the rulers are in unnormalized form.)
Proposition 6 (Sum-of-lengths under combining).
Combining a perfect $(L_{1},M)$-DTS with sum-of-lengths $S_{1}$ and a perfect $(L_{2},M)$-DTS with sum-of-lengths $S_{2}$ as in Theorem 5 yields a perfect $(L_{1}L_{2}M(M+1)+L_{1}+L_{2},M)$-DTS with sum-of-lengths $(L_{2}M(M+1)+1)^{2}S_{1}+S_{2}$.
Proof.
Trivially, $\mathcal{Y}$ and $(L_{2}M(M+1)+1)\mathcal{X}$ contribute $S_{2}$ and $(L_{2}M(M+1)+1)S_{1}$ to the sum-of-lengths so it remains to calculate the contribution of $\mathcal{W}$. The difference between the $v$th and $u$th mark of the ruler in $\mathcal{W}$ corresponding to a fixed $i\in[L_{1}]$, $j\in[L_{2}]$, and $k\in[M(M+1)]$ can be expressed as
$(L_{2}M(M+1)+1)\left(x_{v}^{(i)}-x_{u}^{(i)}\right)+\left(y_{\rho_{k}(v)}^{(j)}-y_{\rho_{k}(u)}^{(j)}\right)\text{.}$
Since the term $y_{\rho_{k}(v)}^{(j)}-y_{\rho_{k}(u)}^{(j)}$ satisfies
$-\frac{L_{2}M(M+1)}{2}\leq y_{\rho_{k}(v)}^{(j)}-y_{\rho_{k}(u)}^{(j)}\leq\frac{L_{2}M(M+1)}{2}\text{,}$
the scale factor $(L_{2}M(M+1)+1)$ guarantees that the ordering of the differences $x_{v}^{(i)}-x_{u}^{(i)}$ is preserved. Assuming without loss of generality that the rulers of $\mathcal{X}$ are normalized so that the length of the $i$th ruler of $\mathcal{X}$ is the difference corresponding to $v=M$ and $u=0$, the same difference will then yield the length of the corresponding rulers in $\mathcal{W}$ as
$(L_{2}M(M+1)+1)x_{M}^{(i)}+\left(y_{\rho_{k}(M)}^{(j)}-y_{\rho_{k}(0)}^{(j)}\right)\text{.}$ (24)
Lastly, sharp $2$-transitivity guarantees that
$\sum_{k=0}^{M(M+1)-1}y_{\rho_{k}(M)}^{(j)}-y_{\rho_{k}(0)}^{(j)}=0$
in which case summing (24) over all $i\in[L_{1}]$, $j\in[L_{2}]$, and $k\in[M(M+1)]$, yields
$L_{2}M(M+1)(L_{2}M(M+1)+1)S_{1}$
which is added to $S_{2}$ and $(L_{2}M(M+1)+1)S_{1}$ to obtain the desired result. ∎
We will now consider each of the cases of $M\in\{1,2,3,4\}$. These cases are both practically the most relevant for applications to higher-order staircase codes but are also theoretically interesting for the reasons just discussed.
VI-A The Case of $M=1$
For $(L,1)$-DTSs, we have the trivial bound
$\max_{\ell\in[L]}d_{1}^{(\ell)}\geq L$
and trivially
$\sum_{\ell=0}^{L-1}d_{1}^{(\ell)}\geq\frac{L(L+1)}{2}$ (25)
with both of these bounds holding with equality for the perfect $(L,1)$-DTS given by $\{(0,1),(0,2),\dots,(0,L)\}$.
VI-B The Case of $M=2$
For $(L,2)$-DTSs, it is known that comparing the parities of (22) and (23) slightly strengthens the trivial bound to
$\max_{\ell\in[L]}d_{2}^{(\ell)}\geq\begin{cases}3L&\text{if }L\equiv 0\text{ or }1\pmod{*}{4}\\ 3L+1&\text{if }L\equiv 2\text{ or }3\pmod{*}{4}\end{cases}\text{.}$ (26)
From this and (22), it straightforwardly follows that
$\sum_{\ell=0}^{L-1}d_{2}^{(\ell)}\\ \geq\begin{dcases}\frac{3L(3L+1)}{4}&\text{if }L\equiv 0\text{ or }1\pmod{*}{4}\\ \frac{(3L-1)3L}{4}+\frac{3L+1}{2}&\text{if }L\equiv 2\text{ or }3\pmod{*}{4}\end{dcases}$ (27)
where the first inequality holds with equality if and only if the set of distances is precisely $\{1,2,\dots,3L\}$ and the second inequality holds with equality if and only if set of distances is precisely $\{1,2,\dots,3L-1,3L+1\}$.
An explicit construction of $(L,2)$-DTSs achieving the bound (26) exists for arbitrary $L$ and was given by Skolem [38] in the case of $L\equiv 0\text{ or }1\pmod{*}{4}$ and by O’Keefe [39] in the case of $L\equiv 2\text{ or }3\pmod{*}{4}$. In the case of $L\equiv 0\text{ or }1\pmod{*}{4}$, by perfectness, the bound (27) is automatically achieved. In the case of $L\equiv 2\text{ or }3\pmod{*}{4}$, the construction of [39] happens to preclude $3L$ from the set of distances and hence also achieves (27). Note that Example 5 demonstrates how it is possible to achieve (26) without achieving (27). Nonetheless, the specific constructions of [38] and [39] simultaneously achieve (26) and (27) for arbitrary $L$ settling this case. We provide the constructions of [38, 39] in the Appendix for the reader’s convenience since [38, 39] consider a different but equivalent problem to the $(L,2)$-DTS problem requiring a well-known but slightly tedious translation.
VI-C The Case of $M=3$
For $(L,3)$-DTSs, the trivial bound is
$\max_{\ell\in[L]}d_{3}^{(\ell)}\geq 6L$ (28)
and it is a longstanding conjecture of Bermond [40] that this is achievable for all $L$ excluding $L\in\{2,3\}$. Theorem 5 implies that there are infinitely many values of $L$ for which (28) is achieved, by, for example, repeatedly combining the perfect $(1,3)$-DTS $\{(0,1,4,6)\}$ with itself. Bermond’s conjecture has been verified for all $L\leq 1000$ in [41] by using a wide variety of recursive constructions to cover gaps as needed.
Proposition 7.
The sum-of-lengths of an $(L,3)$-DTS satisfies
$\sum_{\ell=0}^{L-1}d_{3}^{(\ell)}\geq 5L^{2}+L\text{.}$ (29)
Proof.
By the distinctness of positive differences, we have
$\sum_{u=1}^{6L}u$ $\leq\sum_{\ell=0}^{L-1}\sum_{i=0}^{3}\sum_{j=i+1}^{3}d_{j}^{(\ell)}-d_{i}^{(\ell)}$
$=4\sum_{\ell=0}^{L-1}d_{3}^{(\ell)}-\sum_{\ell=0}^{L-1}(d^{(\ell)}_{1}-d^{(\ell)}_{0})+(d^{(\ell)}_{3}-d^{(\ell)}_{2})$
$\leq 4\sum_{\ell=0}^{L-1}d_{3}^{(\ell)}-\sum_{u=1}^{2L}u$
which yields the desired result after rearranging. ∎
It is not the case that $(L,3)$-DTSs achieving (28) such as those listed in [42] achieve (29). However, we conjecture that it is possible to simultaneously achieve (28) and (29) for all $L$ excluding $L\in\{2,3,4,5\}$. By using a computer-search technique described in Section VII, we found $(L,3)$-DTSs simultaneously achieving (28) and (29) for all $6\leq L\leq 15$ and these are provided in the Appendix. In the cases of $L\in\{2,3,4,5\}$, one or both of (28) and (29) are not achievable but simultaneous absolute minimization of scope and sum-of-lengths is still possible and the optimal solutions are provided in the Appendix. These cases are small enough to be handled by an exact mixed-integer linear programming (MILP) solver by slightly modifying the MILP formulation of the minimum scope DTS problem given in [43]. Alternatively, an exhaustive backtracking DTS search program of Shearer [44, 45] can be used for these small cases.
Next, we consider constructing an infinite family via Theorem 5. For all integers $i\geq 0$, let $\mathcal{Z}_{i}$ be a perfect $(L_{i},3)$-DTS with sum-of-lengths $S_{i}$ obtained by combining $\mathcal{Z}_{i-1}$ and $\mathcal{Z}_{0}$ via Theorem 5 where $\mathcal{Z}_{0}$ is a fixed perfect $(L_{0},3)$-DTS with sum-of-lengths $S_{0}$. Proposition 6 provides linear recurrences for $L_{i}$ and $S_{i}$ which we can solve to get
$L_{i}=\frac{1}{12}\left((12L_{0}+1)^{i+1}-1\right)$
and
$S_{i}=\frac{6S_{0}}{6L_{0}^{2}+L_{0}}L_{i}^{2}+\frac{S_{0}}{6L_{0}^{2}+L_{0}}L_{i}\text{.}$
[…]
$S_{i}=\left(5+\frac{1}{6L_{0}+1}\right)L_{i}^{2}+\left(1-\frac{L_{0}}{6L_{0}+1}\right)L_{i}\text{.}$
Since the largest $L_{0}$ for which we have exhibited a perfect $(L_{0},3)$-DTS achieving (29) in the Appendix is $L_{0}=15$, the best result we can get is as follows:
Proposition 8.
There are infinitely many values of $L$ for which a perfect $(L,3)$-DTS with sum-of-lengths
$\left(5+\frac{1}{91}\right)L^{2}+\left(1-\frac{15}{91}\right)L$
exists.
VI-D The Case of $M=4$
For $(L,4)$-DTSs, it is known that the trivial bound can be slightly strengthened by parity considerations to
$\max_{\ell\in[L]}d_{4}^{(\ell)}\geq\begin{cases}10L&\text{if }L\equiv 0\pmod{*}{2}\\ 10L+1&\text{if }L\equiv 1\pmod{*}{2}\end{cases}\text{.}$ (30)
Proposition 9.
The sum-of-lengths of an $(L,4)$-DTS satisfies
$\sum_{\ell=0}^{L-1}d_{4}^{(\ell)}\\ \geq\begin{dcases}9L^{2}+\frac{3}{2}L&\text{if }L\equiv 0\pmod{*}{2}\\ 9L^{2}+\frac{3}{2}L+\frac{1}{2}&\text{if }L\equiv 1\pmod{*}{2}\end{dcases}\text{.}$ (31)
Proof.
By the distinctness of positive differences, we have
$2\sum_{\ell=0}^{L-1}d_{4}^{(\ell)}=\sum_{\ell=0}^{L-1}\bigg{[}(d^{(\ell)}_{2}-d^{(\ell)}_{0})+(d^{(\ell)}_{4}-d^{(\ell)}_{2})+(d^{(\ell)}_{1}-d^{(\ell)}_{0})+\\ (d^{(\ell)}_{2}-d^{(\ell)}_{1})+(d^{(\ell)}_{3}-d^{(\ell)}_{2})+(d^{(\ell)}_{4}-d^{(\ell)}_{3})\bigg{]}\geq\sum_{u=1}^{6L}u$
which yields the result after integrality considerations. ∎
We conjecture that both (30) and (31) are simultaneously achievable for all $L$ excluding $L\in\{2,3,4\}$. We provide in the Appendix DTSs simultaneously achieving (30) and (31) for all $5\leq L\leq 10$ excluding $L=9$ for which a search is ongoing. The problem of simultaneously achieving (30) and (31) was considered implicitly by Laufer who solved it in the case of $L\in\{6,8\}$ in [46]. This was done because the structure implied by achieving (31) can be exploited to simplify the enumeration of solutions. A solution achieving both of (30) and (31) for $L=10$ was found by Khansefid et al. [47, 42, 48]. In the cases of $L\in\{5,7\}$, we found solutions with Shearer’s search programs [44, 45]. For $L\in\{2,3,4\}$, it is not only the case that one or both of (30) and (31) is not achievable, but also that we get a tradeoff between minimizing scope and minimizing sum-of-lengths when $L=4$. In this case, we provide a pair of Pareto optimal solutions. These smaller cases again can be handled by MILP or Shearer’s search programs [44, 45].
For all integers $i\geq 0$, let $\mathcal{Z}_{i}$ be a perfect $(L_{i},4)$-DTS with sum-of-lengths $S_{i}$ obtained by combining $\mathcal{Z}_{i-1}$ and $\mathcal{Z}_{0}$ via Theorem 5 where $\mathcal{Z}_{0}$ is a fixed perfect $(L_{0},4)$-DTS with sum-of-lengths $S_{0}$. Solving the linear recurrences for $L_{i}$ and $S_{i}$, we get
$L_{i}=\frac{1}{20}\left((20L_{0}+1)^{i+1}-1\right)$
and
$S_{i}=\frac{10S_{0}}{10L_{0}^{2}+L_{0}}L_{i}^{2}+\frac{S_{0}}{10L_{0}^{2}+L_{0}}L_{i}\text{.}$
[…]
$S_{i}=\left(9+\frac{6}{10L_{0}+1}\right)L_{i}^{2}+\left(\frac{3}{2}-\frac{6L_{0}}{10L_{0}+1}\right)L_{i}\text{.}$
Taking $\mathcal{Z}_{0}$ to be the perfect $(10,4)$-DTS achieving (31) in the Appendix, we get the following:
Proposition 10.
There are infinitely many values of $L$ for which a perfect $(L,4)$-DTS with sum-of-lengths
$\left(9+\frac{6}{101}\right)L^{2}+\left(\frac{3}{2}-\frac{60}{101}\right)L$
exists.
VI-E Implications for Higher-Order Staircase Codes
We now consider some implications of the results of this section for higher-order staircase codes:
VI-E1 Recovery of Tiled Diagonal Zipper Codes
Consider the higher-order staircase code defined by the $(2,S/L)$-nets from Example 1 or Example 2 and the scope and sum-of-lengths optimal $(L,1)$-DTS given by $d_{1}^{(\ell)}=L-\ell$ for $\ell\in[L]$. We get that
@@ VI-E1 Recovery of Tiled Diagonal Zipper Codes
belong to some component code $\mathcal{C}$ for all $n\in L\mathbb{Z}$. This family is precisely the tiled diagonal zipper codes of [11].
VI-E2 Memory Reduction Limits
Tiled diagonal zipper codes are often colloquially described as providing encoding and decoding memory reductions by up to $1/2$ relative to classical staircase codes [2] which are the $L=1$ case for this family. The factor of $1/2$ arises as a certain reasonable limit representing the approximate memory reduction at high code rates. We can compute analogous limits for higher-order staircase codes. In particular, we divide the encoding memory (19) by the encoding memory in the $L=1$ case and take the limit as $S\to\infty$ with $S/L$ held constant so that $L\to\infty$ in each of the cases of $M\in\{1,2,3,4\}$. We will assume for each $M\in\{1,2,3,4\}$ the existence of infinitely many values of $L$ for which the respective sum-of-lengths lower bounds (25), (27), (29), and (31) are achieved. The resulting limits are $1/2$, $3/4$, $5/6$, and $9/11$ respectively for each $M\in\{1,2,3,4\}$. Note that our premise is only conjectured to be true in the cases of $M\in\{3,4\}$ but Propositions 8 and 10 allow us to prove limits of $5/6+1/546$ and $9/11+6/1111$ in those cases instead. We can also obtain decoding memory reduction limits via (20) similarly as $1/2$, $3/4$, $6/7$, and $10/12$ respectively for each $M\in\{1,2,3,4\}$.
VI-E3 Recovery of Self-Orthogonal Convolutional Codes